Which of the following is considered a supply shock?

 45

 Which of the following is considered a supply shock?

 A. an increase in input costs

 B. a decrease in investment

 C. an increase in consumer spending

 D.a decrease in imports


 46

 An economy's output grows by 2.5% annually. Calculate the number of years it takes for the economy's output to double. (Do not use the

 rule of 70. Calculate the exact years, round it to 2 decimal places)

 A.28.75

 B.28

 C.28.07

 D.28.25

 E. 28.40


 47

 If the marginal propensity to consume (MPC) is equal to 0.7, govermment increases spending by SX, and the GDP increases by $1000.

 Calculate SX.
Answer #1

45) answer :- option(A)

Due to the increase in input cost there will be decrease in supply. So it is considered as supply shock.

46)

Compounding rate.

P(t) =P0 (1+(i/100))^ny

If doubles means P(t) =2, P0 =1, i=2.5

ny is number of years

2=1(1+(2.5/100))^ny

2=(1.025)^ny

ny=28.07

Option(C).